# Optical properties and electronic correlations in LaNiO bilayer nickelates under high pressure

### v2

@@ Abstract
We explore the optical properties of LaNiO bilayer nickelates by using density functional theory including a Coulomb repulsion term. Convincing agreement with recent experimental ambient-pressure spectra is achieved for  eV, which permits tracing the microscopic origin of the characteristic features. Simultaneous consistency with angle-resolved photoemission spectroscopy and x-ray diffraction suggests the notion of rather moderate electronic correlations in this novel high- superconductor. Oxygen vacancies form predominantly at the inner apical sites and renormalize the optical spectrum quantitatively, while the released electrons are largely accommodated by a defect state. We show that the structural transition occurring under high pressure coincides with a significant enhancement of the Drude weight and a reduction of the out-of-plane interband contribution that act as a fingerprint of the emerging hole pocket. Finally, we investigate the role of the 2-2 versus 1-3 layer stacking and compare the bilayer nickelate to LaNiO, LaNiO, and NdNiO, unveiling general trends in the optical spectrum as a function of the formal Ni valence in Ruddlesden-Popper versus reduced Ruddlesden-Popper nickelates.
@@ I Introduction
The recent observation of superconductivity with  K in pressurized LaNiO [1, 2, 3] suggested the bilayer Ruddlesden-Popper compounds as an intriguing new member of the steadily growing family of superconducting nickelates [4, 5, 6, 7, 8, 9, 10] and instantly sparked considerable interest [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Despite these efforts, several aspects remain unclear so far. Specifically, the pairing mechanism and its relation to the pressure-driven structural transition from a to an  [1, 23, 29] or space group [31, 37], both involving a suppression of the NiO octahedral rotations, is still intensely debated [12, 13, 14, 28, 22, 23, 39].
The key ingredient to understand superconductivity on a fundamental level is an appropriate electronic structure. This necessitates a careful assessment of ab initio results and specifically raises a question about the role of electronic correlations in LaNiO. Angle-resolved photoemission spectroscopy (ARPES) reported a two-band Fermi surface at ambient pressure [34]. Simultaneously, the formation of an additional Ni--derived flat band around the zone corner,  meV below the Fermi level, has been observed [34]. Density functional theory analysis of the Drude peak in the in-plane optical conductivity similar to Qazilbash et al. [41] suggested that LaNiO features strong electronic correlations, placing it in terms of Mottness close to the reference superconductor LaCuO [30].
@@
This motivated us to explore the optical properties of LaNiO bilayer nickelates from first principles including a Coulomb repulsion term. Convincing agreement with the experimental spectrum at ambient pressure is achieved for  eV, which puts us in position to trace the microscopic origin of its characteristic features. In addition to the in-plane optical conductivity, we also predict the out-of-plane component, uncovering an unexpectedly strong anisotropy that reverses as a function of frequency. Simultaneously, Ni energies consistent with recent ARPES results and accurate lattice parameters are obtained, establishing the notion of rather moderate electronic correlations in LaNiO. Moreover, we provide trends in the optical spectrum due to explicit oxygen vacancies, which predominantly occur at the inner apical sites. The released electrons are largely accommodated by an emergent defect state, in sharp contrast to doping the system.
Subsequently, we predict that the structural transition occurring under high pressure is accompanied by a significant enhancement of the Drude peak and a reduction of the out-of-plane interband contribution, which opens a route to track the proposed changes in Fermi surface topology [1, 11, 12, 13, 14, 28, 23] in future -axis measurements. Finally, we uncover a distinct optical signature of the 2-2 versus 1-3 layer stacking [42, 43] and discuss the bilayer compound in the broader context of related Ruddlesden-Popper and reduced Ruddlesden-Popper nickelates, identifying fundamental differences between the two families, but also general trends in the optical response relating to the formal Ni valence.
II Results and Discussion
@@ II.1 Optical spectrum, dipole transition analysis, and impact of oxygen vacancies
We begin with a detailed analysis of the in-plane optical conductivity of LaNiO at ambient pressure [Fig. 1(a); space group] and close comparison with recently measured data obtained from reflectivity experiments [30]. At least four characteristic interband peaks can be identified in the experimental spectrum: A small prepeak at around  eV (A), two dominant peaks at  eV (B) and  eV (C), and a smaller peak at  eV (D). Interestingly, we find that the simulated spectrum obtained for  eV nicely reproduces these features both qualitatively and quantitatively. The quantitative agreement (i.e., the peak intensities and peak energies) is particularly good for peaks B and D, whereas the energy of peak C is overestimated by  eV. Furthermore, close inspection uncovers a small fifth peak at  eV (E) that is reproduced at  eV. The simulation even predicts a finite spectral weight near peak A. Since the experimental in-plane Drude peak is significantly renormalized [30], such low-energy interband features can be clearly resolved.
@@
The convincing agreement of the simulated spectrum for LaNiO and experimental observations puts us in a position to trace the microscopic origin of its characteristic features. The energy- and -resolved TDMs in conjunction with the orbital-resolved band structure [Fig. 1(f)] allow us to identify spectral peaks B and C to in-plane polarized transitions between Ni and Ni states. The low-energy spectral weight (A) stems from excitations in the immediate vicinity of the Fermi level, particular at finite , which are promoted by the substantial octahedral rotations. At higher energies, transitions from the O valence states to the Ni states can be observed (D, E).
@@
The predicted out-of-plane optical conductivity [Fig. 1(b)], while smaller for low energies as expected in a layered system (see inset), exceeds the in-plane components substantially for  eV and uncovers a strong anisotropy of the optical response. The spectrum features a very pronounced peak at  eV that consists of at least two distinct contributions (A, B), a smaller peak at  eV (C), and a broad peak around  eV (D). The Drude peak is rather negligible. We attribute these features to strong transitions between the occupied âbondingâ Ni states and empty âantibondingâ Ni states (A, B) as well as unoccupied Ni âNi âO hybrid states around the  point (C), but also to excitations from the O valence states to Ni states (D) [Fig. 1(f)]. Interestingly, comparison of the TDMs thus reveals that light with varying polarization excites transitions between highly distinct electronic states in these novel nickelate compounds.
@@
Figure 1(e) visualizes the -point wave functions of selected states in LaNiO that contribute prominently to the optical spectrum. We see that the states extend over the entire bilayer due to the substantial Ni -O hybridization. In the Ni--derived states in the valence band, only the basal oxygen ions are involved. In stark contrast, the Ni--derived states in the valence and conduction band show a strong hybridization with the apical oxygen ions, but simultaneously minor contributions from the basal oxygen ions. Notably, the charge density at the inner apical oxygen site is considerably higher for the final than for the initial states; therefore, even an excitation with out-of-plane polarization is associated with a certain degree of charge transfer. This is consistent with the band structure [Fig. 1(f)], which uncovers an increasing involvement of oxygen in the Ni states with increasing energy.
@@
A further interesting unknown at the present time is the role of oxygen vacancies (Vâs) in the physics of bilayer nickelates [28, 31, 38]. It has been reported that the normal-state properties of LaNiO depend sensitively on the oxygen content; in particular, oxygen appears to control the metallicity of the samples [44, 45, 46]. In order to obtain a first-principles impression of the trends in the optical spectrum, we consider an explicit single V in the unit cell. Surprisingly, we identify inner apical oxygen [site 3; see Fig. 1(d)] as the lowest formation energy site ( eVV). Still, this formation energy is higher than e.g. in perovskite LaNiO ( eV [47]). The formation energies of basal Vâs (site 2,  eVV) and outer apical Vâs (site 1,  eVV) are even further enhanced and reach values reminiscent of SrTiO [48, 49], which suppresses these defects exponentially.
@@
Figure 1(a) shows that the presence of an explicit V leaves the overall structure of the in-plane optical spectrum invariant, but enhances the spectral weight below  eV and reduces the intensity of the two main peaks B and C, rendering values closer to experiment. Furthermore, the right slope of peak C becomes less steep and the spectral weight beyond  eV is increased, which also brings the simulated curve closer to experiment. Simultaneously, peak D can no longer be resolved, which indicates that the present V concentration (corresponding formally to LaNiO) is higher than in the experimental sample. For out-of-plane polarization [Fig. 1(b)], the intensity of peak B is significantly reduced, while peaks C and D are broadened and shifted to higher energies.
@@
Inspection of Fig. 1(g) shows that the overall band structure and the TDMs stay similar to the stoichiometric case, particularly in the ideal bilayer without V. Interestingly, two defect states with strong Ni character emerge at and  eV. One of these states is split off from the conduction band and accommodates the two released electrons. Slight variations of are due to states with Ni character around the  point, which experience a lifting of the degeneracy along the Brillouin zone boundary in the defective bilayer.
@@
Due to this mechanism, the impact of Vâs is largely confined to the defective bilayers, while the remaining system shows an electronic structure close to the stoichiometric compound at only modest electron doping. This is reflected in the renormalizations of the optical spectrum, which we find to be moderate in view of the rather high V concentration considered here. We therefore speculate that the general nesting properties of the stoichiometric compound are relatively robust and may carry over to the Fermi surface in slightly oxygen-deficient samples, resulting in a similar superconducting pairing.
@@ II.2 Correlation dependence of the optical spectrum
Next, we investigate how the ambient-pressure optical spectrum of LaNiO varies with the Coulomb repulsion parameter, and estimate the strength of the electronic correlations in this system by comparison to recent experimental data.
Figure 2(a) shows that has a very strong impact on the in-plane optical conductivity, affecting both its intensity and shape. Specifically, we observe a reduction in overall intensity from to  eV, probably due to decreasing TDMs, and a simultaneously raising energy of peak C due to an increasing separation of the Ni and Ni states (see Supplemental Information). For higher values, the energy of peak B is concomitantly decreased.
We identify the best overall agreement with recent experimental data [30] for  eV [Fig. 2(a)]. While the B-C energy difference is even more accurately described for yet smaller values, the peak intensities are up to twice as high as the experimental curve, whereas the spectral weight for photon energies between - eV is considerably underestimated. Good quantitative agreement is also found for  eV, albeit with overestimated B-C energy difference. For  eV, we find a clear disagreement with experiment that particularly manifests in the strongly overestimated B-C energy difference. This poses an upper limit to the static correlation effects in LaNiO.
Additionally, Fig. 2(b) displays the correlation dependence of the out-of-plane component . The characteristic peak around  eV exhibits only a minor energy shift and predominantly becomes narrower with increasing . In sharp contrast, peak C follows a pronounced nonlinear trajectory towards higher energies with increasing , reflecting the continuous lowering of the occupied Ni states and the concomitant raise of the Ni states, particularly the empty ones around the  point (see Supplemental Information). At present, we are not aware of available experimental data for the out-of-plane component of the optical conductivity. Future measurements, particularly for peak C, would provide intriguing additional information about the correlation effects.
@@
Moreover, recent angle-resolved photoemission spectroscopy (ARPES) on LaNiO identified the formation of Ni--derived states  meV below the Fermi level [34]. For  eV ( eV), we observe the Ni states  meV ( meV) below the Fermi energy (see Supplemental Information). This corroborates our conclusion in favor of a rather moderate  eV.
By comparing the Drude weight obtained from experiments to DFT calculations, previous work found that LaNiO features strong electronic correlations that significantly reduce the kinetic energy of the electrons and place the bilayer nickelate in terms of Mottness close to the parent compound of cuprate superconductors LaCuO [30]. The present quantitative analysis of the optical interband transitions, which reflect the relative energies of the active Ni orbitals and their distance to the O states in the valence band, as well as the consistency with ARPES and XRD rather establish the notion of moderate electronic correlations in this novel high- superconductor. This aligns the bilayer compounds with related nickelate systems [50, 51, 52, 53, 54, 55, 48, 56, 57, 58].

### v3

# Optical properties and electronic correlations in La3Ni2O7 bilayer nickelates under high pressure

@@ Abstract
We explore the optical properties of La3Ni2O7 bilayer nickelates by using density functional theory including a Coulomb repulsion term. Convincing agreement with recent experimental ambient-pressure spectra is achieved for eV, which permits tracing the microscopic origin of the characteristic features. Simultaneous consistency with angle-resolved photoemission spectroscopy and x-ray diffraction suggests the notion of rather moderate electronic correlations in this novel high- superconductor. Oxygen vacancies form predominantly at the inner apical sites and renormalize the optical spectrum quantitatively, while the released electrons are largely accommodated by a defect state. We show that the structural transition occurring under high pressure coincides with a significant enhancement of the Drude weight and a reduction of the out-of-plane interband contribution that act as a fingerprint of the emerging hole pocket. We further calculate the optical spectra for various possible magnetic phases including spin-density waves and discuss the results in the context of experiment. Finally, we investigate the role of the 2-2 versus 1-3 layer stacking and compare the bilayer nickelate to La4Ni3O10, La3Ni2O6, and NdNiO2, unveiling general trends in the optical spectrum as a function of the formal Ni valence in Ruddlesden-Popper versus reduced Ruddlesden-Popper nickelates.
@@ I Introduction
The recent observation of superconductivity with K in pressurized La3Ni2O7 [1, 2, 3] suggested the bilayer Ruddlesden-Popper compounds as an intriguing new member of the steadily growing family of superconducting nickelates [4, 5, 6, 7, 8, 9, 10] and instantly sparked considerable interest [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Despite these efforts, several aspects remain unclear so far. Specifically, the pairing mechanism and its relation to the pressure-driven structural transition from a to an [1, 23, 29] or space group [31, 37], both involving a suppression of the NiO6 octahedral rotations, is still intensely debated [12, 13, 14, 28, 22, 23, 39].
The key ingredient to understand superconductivity on a fundamental level is an appropriate electronic structure. This necessitates a careful assessment of ab initio results and specifically raises a question about the role of electronic correlations in La3Ni2O7. Angle-resolved photoemission spectroscopy (ARPES) reported a two-band Fermi surface at ambient pressure [34]. Simultaneously, the formation of an additional Ni--derived flat band around the zone corner, meV below the Fermi level, has been observed [34]. Density functional theory analysis of the Drude peak in the in-plane optical conductivity similar to Qazilbash et al. [41] suggested that La3Ni2O7 features strong electronic correlations, placing it in terms of Mottness close to the reference superconductor La2CuO4 [30].
@@
This motivated us to explore the optical properties of La3Ni2O7 bilayer nickelates from first principles including a Coulomb repulsion term. Convincing agreement with the experimental spectrum at ambient pressure is achieved for eV, which puts us in position to trace the microscopic origin of its characteristic features. In addition to the in-plane optical conductivity, we also predict the out-of-plane component, uncovering an unexpectedly strong anisotropy that reverses as a function of frequency. Simultaneously, Ni energies consistent with recent ARPES results and accurate lattice parameters are obtained, establishing the notion of rather moderate electronic correlations in La3Ni2O7. Moreover, we provide trends in the optical spectrum due to explicit oxygen vacancies, which predominantly occur at the inner apical sites. The released electrons are largely accommodated by an emergent defect state, in sharp contrast to doping the system.
Subsequently, we predict that the structural transition occurring under high pressure is accompanied by a significant enhancement of the Drude peak and a reduction of the out-of-plane interband contribution, which opens a route to track the proposed changes in Fermi surface topology [1, 11, 12, 13, 14, 28, 23] in future -axis measurements. The impact of different magnetic phases including spin-density waves is investigated. Finally, we uncover a distinct optical signature of the 2-2 versus 1-3 layer stacking [42, 43] and discuss the bilayer compound in the broader context of related Ruddlesden-Popper and reduced Ruddlesden-Popper nickelates, identifying fundamental differences between the two families, but also general trends in the optical response relating to the formal Ni valence.
II Results
@@ II.1 Optical spectrum, dipole transition analysis, and impact of oxygen vacancies
We begin with a detailed analysis of the in-plane optical conductivity of La3Ni2O7 at ambient pressure [Fig. 1(a); space group] and close comparison with recently measured data obtained from reflectivity experiments [30]. At least four characteristic interband peaks can be identified in the experimental spectrum: A small prepeak at around eV (A), two dominant peaks at eV (B) and eV (C), and a smaller peak at eV (D). Interestingly, we find that the simulated spectrum obtained for eV nicely reproduces these features both qualitatively and quantitatively. The quantitative agreement (i.e., the peak intensities and peak energies) is particularly good for peaks B and D, whereas the energy of peak C is overestimated by eV. Furthermore, close inspection uncovers a small fifth peak at eV (E) that is reproduced at eV. The simulation even predicts a finite spectral weight near peak A. Since the experimental in-plane Drude peak is significantly renormalized [30], as one can see by comparing to our simulated result, such low-energy interband features can be clearly resolved.
@@
The convincing agreement of the simulated spectrum for La3Ni2O7 and experimental observations puts us in a position to trace the microscopic origin of its characteristic features. The energy- and -resolved TDMs in conjunction with the orbital-resolved band structure [Fig. 1(f)] allow us to identify spectral peaks B and C to in-plane polarized transitions between Ni and Ni states. The low-energy spectral weight (A) stems from excitations in the immediate vicinity of the Fermi level, particular at finite , which are promoted by the substantial octahedral rotations. At higher energies, transitions from the O valence states to the Ni states can be observed (D, E).
@@
The predicted out-of-plane optical conductivity [Fig. 1(b)], while smaller for low energies as expected in a layered system (see inset), exceeds the in-plane components substantially for eV and uncovers a strong anisotropy of the optical response. The spectrum features a very pronounced peak at eV that consists of at least two distinct contributions (A, B), a smaller peak at eV (C), and a broad peak around eV (D). The Drude peak is rather negligible. We attribute these features to strong transitions between the occupied ’bonding’ Ni states and empty ’antibonding’ Ni states (A, B) as well as unoccupied Ni –Ni –O hybrid states around the point (C), but also to excitations from the O valence states to Ni states (D) [Fig. 1(f)]. Interestingly, comparison of the TDMs thus reveals that light with varying polarization excites transitions between highly distinct electronic states in these novel nickelate compounds.
@@
Figure 1(e) visualizes the -point wave functions of selected states in La3Ni2O7 that contribute prominently to the optical spectrum. We see that the states extend over the entire bilayer due to the substantial Ni -O hybridization. In the Ni--derived states in the valence band, only the basal oxygen ions are involved. In stark contrast, the Ni--derived states in the valence and conduction band show a strong hybridization with the apical oxygen ions, but simultaneously minor contributions from the basal oxygen ions. Notably, the charge density at the inner apical oxygen site is considerably higher for the final than for the initial states; therefore, even an excitation with out-of-plane polarization is associated with a certain degree of charge transfer. This is consistent with the band structure [Fig. 1(f)], which uncovers an increasing involvement of oxygen in the Ni states with increasing energy.
@@
A further interesting unknown at the present time is the role of oxygen vacancies (V’s) in the physics of bilayer nickelates [28, 31, 38]. It has been reported that the normal-state properties of La3Ni2O7 depend sensitively on the oxygen content; in particular, oxygen appears to control the metallicity of the samples [44, 45, 46]. In order to obtain a first-principles impression of the trends in the optical spectrum, we consider an explicit single V in the unit cell. Surprisingly, we identify inner apical oxygen [site 3; see Fig. 1(d)] as the lowest formation energy site ( eVV), which has been confirmed independently by high-resolution electron ptychography [47]. Still, this formation energy is higher than e.g. in perovskite LaNiO3 ( eV [48]). The formation energies of basal V’s (site 2, eVV) and outer apical V’s (site 1, eVV) are even further enhanced and reach values reminiscent of SrTiO3 [49, 50], which suppresses these defects exponentially.
@@
Figure 1(a) shows that the presence of an explicit V leaves the overall structure of the in-plane optical spectrum invariant, but enhances the spectral weight below eV and reduces the intensity of the two main peaks B and C, rendering values closer to experiment. Furthermore, the right slope of peak C becomes less steep and the spectral weight beyond eV is increased, which also brings the simulated curve closer to experiment. Simultaneously, peak D can no longer be resolved, which indicates that the present V concentration (corresponding formally to La3Ni2O6.75) is higher than in the experimental sample. For out-of-plane polarization [Fig. 1(b)], the intensity of peak B is significantly reduced, while peaks C and D are broadened and shifted to higher energies.
@@
Inspection of Fig. 1(g) shows that the overall band structure and the TDMs stay similar to the stoichiometric case, particularly in the ideal bilayer without V. Interestingly, two defect states with strong Ni character emerge at and eV. One of these states is split off from the conduction band and accommodates the two released electrons. Slight variations of can be attributed to states with Ni character around the point, which experience a lifting of the degeneracy along the Brillouin zone boundary in the defective bilayer. Concomitantly, the Fermi surfaces show a strongly reconstructed shape and reduced Fermi velocity (i.e., increased resistivity; blue colors) in the defective bilayer, particularly near the point [Fig. 1(g)].
@@
This mechanism confines the impact of V’s largely to the defective bilayers, while the remaining system shows an electronic structure close to the stoichiometric compound at only modest electron doping. This is reflected in the renormalizations of the optical spectrum, which we find to be moderate in view of the rather high V concentration considered here. We therefore speculate that the general nesting properties of the stoichiometric compound are relatively robust and may carry over to the Fermi surface in slightly oxygen-deficient samples, resulting in a similar superconducting pairing.
@@ II.2 Correlation dependence of the optical spectrum
Next, we investigate how the ambient-pressure optical spectrum of La3Ni2O7 varies with the Coulomb repulsion parameter, and estimate the strength of the electronic correlations in this system by comparison to recent experimental data.
Figure 2(a) shows that has a very strong impact on the in-plane optical conductivity, affecting both its intensity and shape. Specifically, we observe a reduction in overall intensity from to eV, probably due to decreasing TDMs, and a simultaneously raising energy of peak C due to an increasing separation of the Ni and Ni states (see Supplementary Information). For higher values, the energy of peak B is concomitantly decreased.
We identify the best overall agreement with recent experimental data [30] for eV [Fig. 2(a)]. While the B-C energy difference is even more accurately described for yet smaller values, the peak intensities are up to twice as high as the experimental curve, whereas the spectral weight for photon energies between - eV is considerably underestimated. Good quantitative agreement is also found for eV, albeit with overestimated B-C energy difference. For eV, we find a clear disagreement with experiment that particularly manifests in the strongly overestimated B-C energy difference. This poses an upper limit to the static correlation effects in La3Ni2O7.
Additionally, Fig. 2(b) displays the correlation dependence of the out-of-plane component . The characteristic peak around eV exhibits only a minor energy shift and predominantly becomes narrower with increasing . In sharp contrast, peak C follows a pronounced nonlinear trajectory towards higher energies with increasing , reflecting the continuous lowering of the occupied Ni states and the concomitant raise of the Ni states, particularly the empty ones around the point (see Supplementary Information). At present, we are not aware of available experimental data for the out-of-plane component of the optical conductivity. Future measurements, particularly for peak C, would provide intriguing additional information about the correlation effects.
@@
Moreover, recent angle-resolved photoemission spectroscopy (ARPES) on La3Ni2O7 identified the formation of Ni--derived states meV below the Fermi level [34]. For eV ( eV), we observe the Ni states meV ( meV) below the Fermi energy (see Supplementary Information), an important agreement given current debates about the electronic structure of bilayer nickelates. This corroborates our conclusion in favor of eV.
By comparing the Drude weight obtained from experiments to DFT calculations, previous work found that La3Ni2O7 features strong electronic correlations that significantly reduce the kinetic energy of the electrons and place the bilayer nickelate in terms of Mottness close to the parent compound of cuprate superconductors La2CuO4 [30]. The present quantitative analysis of the optical interband transitions, which reflect the relative energies of the active Ni orbitals and their distance to the O states in the valence band, as well as the consistency with ARPES and XRD rather establish the notion of moderate electronic correlations in this novel high- superconductor.
The rare-earth nickelates [51] including the infinite-layer compounds [52, 53, 54, 55] are known to exhibit a more covalent nature and distinct Ni-O hybridization than related cuprates. Consequently, the latter are often described by higher Hubbard- values such as eV [56, 57] than nickelate systems (- eV [58, 59, 60, 61, 62, 63, 64]). It has been reported that the bilayer nickelates show a strong involvement of the oxygen system as well [23, 31]. The resulting more delocalized wave functions rationalize the trend of reduced electronic correlations relative to the paradigmatic cuprate superconductors.